What is the variance of first 50even natural numbers ?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
Its formula is 
(n² -1)/3
Hence variance is = (50²-1)/3 = 833
Regards
Arun (askIITians forum expert)
Khimraj
5 years ago
variance is given by σ =(∑Xi2/N) – (X¯ )2
X¯ = 2+4+6+......+100 
        50
X¯  = 51
(X¯)= 2601
(∑Xi2/N)  = (22+42+.....+1002)/50 = 4(12+22+.......+502)/50 =4(50*51*101/50*6) = 3434
so σ 2= 3434-2601 = 833.
Hope it clears if still u have doubt u can ask me. If u like anser then please approve it.